### v1

@@ Results
Using permutation-based importance testing we observe the expected result that the SICCT test is overwhelmingly the most important indicator of a confirmed breakdown, accounting for around 23% of model accuracy. Figure 5 shows the accuracy contribution and ranking of each model feature. After the SICCT result, the most significant features were the date of testing, the location, the type of testing event (i.e. reason for testing), number of days since the herd was last in breakdown, the size of the herd, the abundance of badgers near the holding, and the APHA risk score assigned to the herd.
Figure 5: The relative importance of model features (risk factors), as tested by permutation-based importance testing, with a random control variable (CONTROL). Features marked ** are from the veterinary dataset that only covers around 10% of the full dataset, and so likely have underestimated importance. Features marked * refer to previous tests or breakdowns and will be left-censored where the previous test or breakdown is before the first date in the dataset. Error bars show 95% confidence interval.
Of those features that can be considered direct risk factors the most important were (in order of importance): location, time since the last breakdown, herd size, badger abundance, movements into the herd over the year prior to the test, the result and time of the previous testing event, and the veterinary practice conducting the test. Animal movements were important, with all types of movement contributing more than the random control variable, but movements into the herd over one year was the most significant by some margin; other periods and movements out of herd were substantially less significant. Movements from farms with breakdowns within 2 years (prior or since) appear to make little contribution. The veterinary practice conducting the test was important but the tuberculin batch had very little importance.
Veterinary practice factors
We further interrogated whether the contribution of veterinary practice to the model is correlated with other risk factors in the model. We define “accuracy” as the proportion of times a positive herd-level test leads to a confirmed breakdown and a negative test does not lead to a confirmed breakdown. Most veterinary practices are accurate more often than not by this definition, but accuracy is substantially lower for some (60% or lower, Figure 6a). This variation is not easily explained by random chance; in a binomial test where we assume the average accuracy per herd test for all practices, 58.6% of practices have accuracy outside the expected range (cutoff $p\geq 0.05$). One factor that influences this is the mean herd size tested by the veterinary practice (Figure 6b). Mean herd size tested by the practice is correlated with accuracy with Pearson $r=-0.41$ ($p\approx 10^{-17}$).
(a)Refer to caption (b)Refer to caption
Figure 6: (a) Accuracy, defined as the proportion of times a positive herd-level test leads to a confirmed breakdown and a negative test does not lead to a confirmed breakdown, by veterinary practice. (b) Correlation between veterinary practice accuracy and the mean herd size tested by the practice. Pearson r=-0.41.
The model can be shown to correct for these factors by increasing accuracy for almost all practices (Figure 7a). Practice accuracy is increased by greater proportions for practices that test larger mean herd sizes (Figure 7b).
(a)Refer to caption (b)Refer to caption
Figure 7: (a) The accuracy increase provided by the diagnostic model, by the number of veterinary practices. (b) The accuracy increase achieved in practices by the mean herd size tested by the practice.
@@ Discussion
The usual interpretation (threshold on measurement) of the standard SICCT test is tuned so as to maintain an acceptably low level of false positive results, i.e. to minimise the number of times where herd level restrictions are put into place unnecessarily, where there is no underlying true infections in the herd. This is sensible both because the relatively low overall level of infection across GB (as of Dec 2023, in England a herd prevalence of 4.6% [20]) means that it is mostly truly negative cattle that are being tested, and because the implications of a positive test are severe for both the keeper and the government (i.e. herd restrictions and multiple follow-up tests). While the SICCT test is interpreted in the same fashion across all OTF herds, the epidemiological context is already taken into account in that the frequency of routine bTB herd surveillance testing varies with the regional risk of bTB, and for the vast majority of herds where there is already a strong suspicion of infection—bTB breakdown herds that have had their OTF status suspended due to detection of at least one SICCT test reactor—a more severe interpretation of the SICCT test is automatically used. In this case a larger likelihood of false positive test results is deemed acceptable in exchange for an enhanced probability of removing all infected animals from the herd. This work extends this concept by evaluating whether an assessment of epidemiological risk (i.e. a prior probability that a herd undergoing testing is more or less likely to already harbour infected cattle) could be used to alter the interpretation of the SICCT test, even when the herd in question does not already have its OTF status suspended or withdrawn at the point of testing. That is, even when the herd remains OTF after a clear SICCT performed at the standard interpretation (e.g. where only inconclusive reactors are detected), or in those breakdown herds with OTF status suspended in the Low Risk Area of England and in Scotland where the severe interpretation of the SICCT test is not automatically applied. This concept has already been exploited to a more limited degree in Scotland [21] and re-evaluated in England to show proof of concept [22].
However, those previous results have not considered the full implications that result from either catching some infected herds early or reducing restrictions on others. A machine learning based approach already used to evaluate GB testing [9] provides a platform on which to do a more sophisticated analysis. To accomplish this we evaluated test results under a specific set of circumstances where subsequent testing is likely to be a reliable indicator of the true infection status of a herd. Specifically, we can evaluate where a historical SICCT test should be viewed differently by looking ahead at whether a confirmed breakdown occurred or not.. If a clear or inconclusive test is followed soon after by a positive test with a confirmed reactor, this is a reasonable indicator that there may have been some benefit to the original clear test having been interpreted more severely. Similarly, if an inconclusive test is followed up soon after by a clear set of follow-up tests, this is an indicator that the herd could have safely been relieved of further restrictions beyond the initial culling of test positive animals, thereby substantially reducing the burden to both the farmer, and the veterinary testing services.
Whether or not machine-learning augmentation of testing could be used in practice depends on many factors outside this model including the regulatory framework; we show that in principle such an approach has potential merit and warrants further investigation. By testing this proposition across the entire record of herd tests using well-established machine learning techniques, we show that simple indicators of epidemiological risks can result in both substantial numbers of herds where an incipient infection can be identified early, as well as many herds that could have been saved a period of restriction and further testing. This approach could potentially be broadened to any disease diagnostic situation where sufficient data is available and could be particularly useful for non-statutory diseases where there is a substantial need for improved testing and greater flexibility in farmer-led management of the disease, as with bovine viral diarrhoea and Johne’s disease.
Our simulation model results, which consider more extensively the transmission dynamics of the bacteria beyond the individual herd, shows that increasing the HSe has a variable impact on the total number of breakdowns, even though there is an impact in reducing the transmissions between farms, and a notable decrease in the number of individual reactors. When HSp is increased while maintaining HSe, we see a decrease in the number of breakdowns and individual reactors in both simulated areas.
@@
Despite the complexity of the risk factors affecting bTB transmission, the key protagonists are well known and are confirmed in our risk factor analysis. However, one unexpected outcome was the role that date of testing played in model accuracy, implying that the weighting of other risk factors is not constant, with some becoming more or less important over time in a changing risk landscape. The month in which a test is taken has far lower importance in the model, implying annual seasonality is less important.
In our analysis of veterinary-related risk factors we see that the veterinary practice conducting the test plays a significant role, despite low coverage of the test dataset with veterinary information. We do not see any risk associated with tuberculin batch, though this may be because the test dataset coverage for tuberculin batch information is very low. The variability in testing outcomes by practice is partially correlated to the average size of herd tested and this may provide some clues as to reasons behind the result. This analysis does not provide evidence to indicate that variability by practice is a result of practice management, as there are many potential confounders; for example, there may be local disease-related risks which are not included in the model. Nonetheless, by showing that the model can identify variability across practices, it does provide a foundation for further interrogation of management factors that could be used to improve outcomes for veterinarians.
@@
We have shown with historical data, that a substantial benefit can be gained by considering complex epidemiological interactions in a single model. Further exploration of sensitivity/specificity benefit trade-offs could also result in improved outcomes. While the underlying analytical model is sophisticated, its outcomes could be easily implemented in real time with software that could be used on a hand held device, and the risk factors displayed graphically in such a way that the results could be useful in the development of a more targeted testing policy, and supplied directly to veterinarians and farmers.
Acknowledgements
This work was supported by the GB bovine TB research budget (grant SE3330) held and administered centrally by Defra on behalf of England, Scotland and Wales.

### v2

@@ Results
Using SHAP importance testing we do not observe any unexpected results, those factors known to he highly associated with bTB risk appear to make significant contributions in our model. Figure 3 shows the importance ranking of each model feature for which the feature’s absolute SHAP values are significantly greater than those of a random feature. The most significant features were the number of animals tested, the APHA risk scoring, volume of movements into the herd in the last four years, date of testing, recency of a breakdown in the herd, herd location, and the result of the previous test.
Figure 3: The relative importance of model features (risk factors), as tested by SHAP importance testing, with a random control variable. Only feature whose absolute SHAP values are significantly greater than the random feature (Mann-Whitney U test, $p<0.01$) are shown. Features marked * refer to previous tests or breakdowns and will be left-censored where the previous test or breakdown is before the first date in the dataset.
Of those features that can be considered direct risk factors the most important were (in decreasing order of importance): number of animals, movements into the herd, recency of last breakdown, location, and previous test result. Herd size and animal movements were important, with all types of movement contributing more than the random control variable, but movements into the herd over the two to four years prior to the test was the most significant by some margin; other periods and movements out of herd were less significant. Movements from farms with breakdowns within two years appear to make only a small contribution. Neither the veterinary practice conducting the test nor the tuberculin batch had significant importance.
@@ Discussion
In this study we have developed a model that predicts the risk landscape associated with bTB breakdowns. We apply this to provide an augmented interpretation of a bTB SICCT test in routine whole-herd testing events. Using a scheme where herd-level specificity is maintained, we are able to improve herd-level sensitivity such that a significant number of false negative test results are caught early in the testing over historical data (Objective 1). By tuning the decision threshold it is also possible to reduce false positives, with that tuning allowing for a balance between the two outcomes. Undetected infections are important because they both potentially extending periods of herd restriction once infection is discovered, and they allow for onward transmission to other herds and to wildlife that would not occur, if detected earlier. On the other hand, because the vast majority of herds do not harbour infection, even small changes in specificity can have a much greater influence on more herds. While the HSe-focussed model identifies 240 herds that eventually had confirmed breakdowns but had negative tests, the HSp-focussed model identifies 5225 herds that had a positive test but did not eventually have a confirmed breakdown (in 2020). As all breakdowns have a substantial impact simply due to the restrictions that are imposed as a result, this trade-off must carefully consider the approach that would have the greatest positive impact, if implemented in practice.
We further considered this trade-off with a simulation model that looked at the short term onward impact that missed breakdowns have at a regional level (i.e. considering on other herds). We further compare two different geographical areas—one in Devon where there has been a decades long bTB problem and the other in Derbyshire, where the history of bTB is more recent. While increasing the HSe notably decreases the number of individual reactors in herds in both areas, and thus there is overall a net reduction in infections, the impact on the overall number of breakdowns shows inconclusive benefits, at least in the short term, suggesting that onward benefits impacting the national epidemic may take longer to become manifest. Nevertheless, the overall reduction in test reactors in all cases suggests that if applied “in the field” in representative areas, this augmented testing method may be successfully used either to reduce impact and burden on farmers and herds. (Objective 2.)
@@
Despite the complexity of the risk factors affecting bTB transmission, the key protagonists are well known and are confirmed in our risk factor analysis. However, one unexpected outcome was the role that date of testing played in model accuracy. This implies that the weighting of other risk factors is not constant, with some becoming more or less important over time in a changing risk landscape. The month in which a test is taken has far lower importance in the model, implying annual seasonality is less important.
We included the veterinary practice and tuberculin batch data because anecdotally these are considered to have some impact on testing outcomes and we wished to test this. Ultimately neither of these features had significant importance in the model. However, the coverage for both batch and veterinary practice were both low; so with improved coverage, either or both could be found to have greater importance (Objective 3.)
The usual interpretation (threshold on measurement) of the standard SICCT test is tuned so as to maintain an acceptably low level of false positive results, i.e. to minimise the number of times where herd level restrictions are put into place unnecessarily, where there is no underlying true infection in the herd. This is sensible both because the relatively low overall level of infection across GB (as of Dec 2023, in England a herd prevalence of 4.6% [20]) means that it is mostly truly negative cattle that are being tested, and because the implications of a positive test are severe for both the keeper and the government (i.e. herd restrictions and multiple follow-up tests). While the SICCT test is interpreted in the same fashion across all OTF herds, the epidemiological context is already taken into account in that the frequency of routine bTB herd surveillance testing varies with the regional risk of bTB, and both for where there is already a strong suspicion of infection—bTB breakdown herds that have had their OTF status removed (either suspended/OTFS or withdrawn/OTFW) or for cattled traced from breakdown herds—a more severe interpretation of the SICCT test is automatically used. In this case a larger likelihood of false positive test results is deemed acceptable in exchange for an enhanced probability of removing all infected animals from the herd. Our analysis extends this concept by evaluating whether an assessment of epidemiological risk (i.e. a prior probability that a herd undergoing testing is more or less likely to already harbour infected cattle) could be used to alter the interpretation of the SICCT test, even when the herd in question does not already have its OTF status suspended or withdrawn at the point of testing. That is, even when the herd remains OTF after a clear SICCT performed at the standard interpretation (e.g. where only inconclusive reactors are detected), or in those breakdown herds with OTF status suspended in the Low Risk Area of England and in Scotland where the severe interpretation of the SICCT test is not automatically applied. This concept has already been exploited to a more limited degree in Scotland [21] and re-evaluated in England to show proof of concept [22].
However, those previous results have not considered the full implications that result from either catching some infected herds early or reducing restrictions on others. A machine learning based approach already used to evaluate GB testing [9] provides a platform on which to do a more sophisticated analysis. To accomplish this, we evaluated test results under a specific set of circumstances where subsequent testing is likely to be a reliable indicator of the true infection status of a herd. Specifically, we can evaluate where a historical SICCT test should be viewed differently by looking ahead at whether a confirmed breakdown occurred or not. If a clear or inconclusive test is followed soon after by a positive test with a confirmed reactor, this is a reasonable indicator that there may have been some benefit to the original clear test having been interpreted more severely. Similarly, if an inconclusive test is followed up soon after by a clear set of follow-up tests, this is an indicator that the herd could have safely been relieved of further restrictions beyond the initial culling of test-positive animals, thereby substantially reducing the burden to both the farmer, and the veterinary testing services.
Whether machine-learning augmentation of testing could be used in practice depends on many factors outside this model including the regulatory framework; we show that in principle such an approach has potential merit and warrants further investigation. By testing this proposition across the entire record of herd tests using well-established machine learning techniques, we show that simple indicators of epidemiological risks can result in both substantial numbers of herds where an incipient infection can be identified early, as well as many herds that could have been saved a period of restriction and further testing. This approach could potentially be broadened to any disease diagnostic situation where sufficient data is available and could be particularly useful for non-statutory diseases where there is a substantial need for improved testing and greater flexibility in farmer-led management of the disease, as with bovine viral diarrhoea and Johne’s disease.
@@
Conclusion
We have shown with historical data that a substantial benefit can be gained by considering complex epidemiological interactions in a single model. Further exploration of sensitivity/specificity benefit trade-offs could also result in improved outcomes. While the underlying analytical model is sophisticated, its outcomes could be easily implemented in real time with software held on a hand held device, and the risk factors displayed graphically in such a way that the results could be useful in the development of a more targeted testing policy, and supplied directly to veterinarians and farmers.
Dedication
In memoriam Thomas Patrick Doherty 1975–2024
The simulation model section of this paper could not have been achieved without the groundwork laid by Tom, the key developer of the simulation model used in this paper.
Tom was a lynch-pin member of our team and the most hard-working, dedicated, kind, and helpful gentleman one could ever hope to meet. Tom sadly passed away before the conclusion of this paper, but his work will live on and continue to help us. He will be sorely missed by all of us. Rest in peace, Tom.
Funding sources
This work was supported by the GB bovine TB research budget (grant SE3330, awarded to RRK) held and administered centrally by Defra on behalf of England, Scotland and Wales, and also by the Roslin Institute ISP2(theme 3), BBSRC (grant BBS/E/D/20002174, awarded to Roslin Institute, funds RRK and CJB). Defra approved release of the manuscript, but otherwise the funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
Data Availability Statement
Data for this analysis were extracted from the Animal and Plant Health Agency (APHA) bTB surveillance database (SAM) [12] and the Cattle Tracing System (CTS) [13] databases. These data were made available under a Data Sharing Agreement between APHA and the University of Edinburgh. Data on the veterinary practice conducting the test and the tuberculin batches used were supplied by UK Farmcare [14]. Data requests may be addressed to enquiries@apha.gov.uk for APHA, and admin@ukfarmcare.com for UKFarmcare.
License
For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.